# La RoSA: Enhancing LLM Efficiency via Layerwise Rotated Sparse Activation

### v1

@@ 4.2 Layerwise Orthogonal Rotation
The limitations of magnitude pruning drive us to explore a more effective approach for identifying activations that are inherently easier to prune. Inspired by rotation-based quantization methods (Ashkboos et al., 2024b), we attempt to obtain a rotated activation through orthogonal transformations, after which the channel importance can be easily distinguished. We propose using an orthogonal matrix ${\mathbf{Q}}_{l}$ to rotate and prune the input activation (i.e., ${\bm{h}}_{1}$ and ${\bm{h}}_{3}$) of $l$-th layer, and reverse the rotation using $\mathbf{Q}^{\top}_{l}$ after pruning. To construct ${\mathbf{Q}}_{l}$, we use PCA to leverage it’s inherent ability to identify principal components. Specifically, we select a calibration dataset consist of $M$ sequences, feed it into the model and forward pass through each layer. Let ${\mathbf{X}}_{l}^{i}$ denotes the input activation of layer $l$ for $i$-th sequence. We compute the covariance matrix ${\mathbf{X}}_{l}^{i}({\mathbf{X}}_{l}^{i})^{T}$ of each ${\mathbf{X}}_{l}^{i}$ and average over all sequences:
@@
It can be observed that each layer must use the same ${\mathbf{Q}}_{l}$ due to residual connections between layers. We initially attempt to use a single global ${\mathbf{Q}}$ for all layers, but quickly find out that the optimal orthogonal matrix for each layer tends to be very different. To address this, a residual adapter ${\mathbf{Q}}_{l}^{T}{\mathbf{Q}}_{l+1}$ is applied to the output of each layer to realize layer-specific rotation, as shown in Figure 2. Although this introduce small additional operations, we demonstrate in Section 5.3 that such operations preserve the optimal rotation for each layer and improve the overall efficiency. In Appendix A, we provide a theoretical analysis of rotation-based activation pruning, demonstrating that the constructed rotation matrix effectively reduces empirical error within each layer, thereby outperforming magnitude pruning methods in accuracy.
@@ 4.3 Introducing Consistent Activation Sparsity
where ${\mathbf{Q}}_{l}^{T}{\mathbf{W}}_{l}^{T}$ can be precomputed and merged into the weights to avoid online computation. Therefore, Equation 9 can be rewritten as :
@@ 4.4 Hardware-efficient Customized Kernels
Introducing activation sparsity alone is insufficient to achieve end-to-end LLM inference acceleration on hardware (Frantar & Alistarh, 2023). To translate the activation sparsity induced by LaRoSA into wall-clock inference speed-up, we need to implement a fast and efficient customized GPU kernel. Building on the Triton-based kernel introduced by DejaVu (Liu et al., 2023) and TEAL (Liu et al., 2024a), we implement a GEMV kernel that provides practical and consistent token generation speed-up. 1) We store weight matrices ${\mathbf{W}}$ in column-major format in memory to facilitate the selective loading of weight columns corresponding to non-zero activations, as memory coalescing could significantly improve memory bandwidth utilization (Dao et al., 2022; Ivanov et al., 2020). 2) We fuse the Top-K function into the matrix-vector multiplication, identifying the indices of the columns in the weights that need to be retained during the forward pass. 3) We selectively load the sparse activations and corresponding weight columns from memory, transfer them to the computation units, perform the multiplication, and store the result.
@@ 5 Experiments
Table 1: Comparison of the average accuracy on seven zero-shot tasks and 5-shot MMLU. The first column indicates the actual model-level sparsity. CATS cannot achieve 50% model-level sparsity as it only sparsifies the MLP’s Down projections. Full results are in  Appendix C.
@@ 5.1 Experimental Setups
Models and Baselines. To test the robustness and scalability of LaRoSA, we conducted experiments using 4 leading LLMs: LLaMA3 (Team, 2024), LLaMA2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), and Qwen2.5 (Qwen et al., 2025). The model size is selected from 7B and 70B parameters. We compare with state-of-the-art (SOTA) baselines, including CATS (Lee et al., 2024) and TEAL (Liu et al., 2024a). For SOTA baselines, we choose not to quote the results from their papers in order to make fair comparisons under the same model-level sparsity setting and to impose full sparsity in the input sequence length dimension.
Implementation Details. We use the WikiText2 train set (Merity et al., 2016) as calibration dataset for LaRoSA and other reproducible works. We randomly select 16 sequences with sequence length of 2048 tokens to compute the rotation matrices ${\mathbf{Q}}$ for LaRoSA and empirical distributions for CATS and TEAL. The computation of ${\mathbf{Q}}$ is performed on 8x80G A100 GPUs, taking approximately 12 minutes to complete for the LLaMA3 70B model. For sparsity coefficient $\alpha$, we employ Grid Search to find the optimal hyperparameter for each model, as shown in Appendix B Algorithm 1. The optimal $\alpha$ for each activation type of models is presented in Appendix B Table 11. For a fair comparison, we implement full activation sparsification at prefill stage for all methods including TEAL.
Evaluation. We report the perplexity (PPL) score on WikiText2 test set for generation task. All models are evaluated on the same 128 random samples with a 2048-token context length. For comprehensive assessment, we utilize lm-evaluation-harness(Gao et al., 2023) to assess LaRoSA performance on a variety of downstream tasks, including 0-shot tasks for ARC-Easy (Clark et al., 2018), ARC Challenge (Clark et al., 2018), PIQA (Bisk et al., 2020), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018) and WinoGrande (Sakaguchi et al., 2021). We also report 5-shot accuracy for MMLU (Hendrycks et al., 2020).
@@
Dense 109.95 (1.00$\times$) 101.25 (1.00$\times$) 22.05 (1.00$\times$) 107.54 (1.00$\times$) 19.32 (1.00$\times$) 105.88 (1.00$\times$)
0% 96.75 (0.88$\times$) 92.13 (0.89$\times$) 20.32 (0.90$\times$) 100.97 (0.90$\times$) 56.33 (0.92$\times$) 56.33 (0.88$\times$)
A100 25% 125.34 (1.14$\times$) 111.37 (1.10$\times$) 24.70 (1.12$\times$) 123.67 (1.15$\times$) 22.60 (1.17$\times$) 118.59 (1.12$\times$)
50% 151.73 (1.38$\times$) 131.62 (1.30$\times$) 29.77 (1.35$\times$) 152.71 (1.42$\times$) 29.17 (1.51$\times$) 148.23 (1.40$\times$)
75% 188.01 (1.72$\times$) 174.15 (1.72$\times$) 38.59 (1.75$\times$) 180.67 (1.68$\times$) 33.42 (1.73$\times$) 186.35 (1.76$\times$)
Dense 72.47 (1.00$\times$) 67.19 (1.00$\times$) 15.75 (1.00$\times$) 71.70 (1.00$\times$) 14.80 (1.00$\times$) 72.89 (1.00$\times$)
0% 65.95 (0.91$\times$) 60.47 (0.90$\times$) 14.33 (0.91$\times$) 63.81 (0.89$\times$) 13.62 (0.92$\times$) 72.89 (0.90$\times$)
H20 25% 85.51 (1.18$\times$) 79.28 (1.18$\times$) 19.21 (1.22$\times$) 83.17 (1.16$\times$) 18.35 (1.24$\times$) 65.60 (1.16$\times$)
50% 104.36 (1.44$\times$) 94.74 (1.41$\times$) 22.84 (1.45$\times$) 98.95 (1.38$\times$) 21.46 (1.45$\times$) 84.55 (1.40$\times$)
75% 136.97 (1.89$\times$) 126.99 (1.89$\times$) 29.92 (1.90$\times$) 132.64 (1.85$\times$) 28.42 (1.92$\times$) 102.05 (1.90$\times$)
@@ 5.2 Main Results
Language Generation Task. First, we evaluate the accuracy of LaRoSA with SOTA baselines on language generation task. As indicated in Table 2, LaRoSA notably outperforms other baselines on perplexity across various models and different sparsity settings. Under a low sparsity setting like 25%, LaRoSA achieves a minimal PPL loss of merely 0.1 on LLaMA3-70B. This significantly surpasses previous methods, such as CATS with a PPL loss of 0.71 and TEAL with a PPL loss of 1.1. LaRoSA maintains its advantage even at high sparsity levels, such as 60%. For instance, on Qwen2.5-72B, LaRoSA incurs only a 1.03 PPL loss, considerably outperforming TEAL, which suffers a loss of 3.87. Moreover, on LLaMA2-7B, LaRoSA at 50% sparsity even outperforms the previous SOTA methods under a 25% sparsity setting, highlighting its superior efficiency. In TEAL’s paper, they only sparsify 99% of tokens in the prefill stage because the empirical thresholds they use struggle to sparsify initial tokens. In contrast, LaRoSA does not suffer this performance degradation and can achieve full sparsity in the input sequence length dimension.
@@
Complex Problem Solving Tasks. To further evaluate the effectiveness and robustness of LaRoSA, we conduct experiments on complex tasks such as MATH500 (Lightman et al., 2024), GPQA-Diamond (Rein et al., 2024), and AIME’24 (AIME, 2025) using two reasoning models: DeepSeek-R1-Distill-Llama3-8B and DeepSeek-R1-Distill-Qwen2.5-7B (DeepSeek-AI, 2025). We utilize the Alpaca dataset, which is not inherently related to complex reasoning tasks, to derive the rotation matrix ${\mathbf{Q}}_{l}$ for each layer. The evaluation of LaRoSA on these complex tasks is carried out with the Hugging Face Open-R1 repository (Face, 2025). As shown in Table 4, our experimental results indicate that at a sparsity level of 25%, both reasoning models exhibit only a slight performance degradation on complex tasks. This finding demonstrates that the LaRoSA method maintains its effectiveness even when utilizing data that is not directly related to complex reasoning tasks for generating the rotation matrix.
@@
Inference Speed Up. We demonstrate consistent end-to-end efficiency improvements with LaRoSA. For this, we collected ten samples, each consisting of 128 tokens, from various test datasets and generated new sequences with lengths ranging from 128 to 2048 tokens. This was done to evaluate performance across different generation lengths. We applied 50% sparsity to LLaMA3-8B and Qwen2.5-7B using optimal sparse patterns obtained from LaRoSA and TEAL. As depicted in Figure 4, our LaRoSA realizes more consistent wall-clock time speed-ups compared to TEAL. This improvement is attributed to the TopK strategy which precisely zeros out activation entries needed for a specific sparsity level. As the output sequence length increases, LaRoSA outperforms TEAL in overall speed-up efficiency, demonstrating its effectiveness and robustness. Moreover, we evaluated the end-to-end token generation speed-up on different types of models at a fixed output sequence length of 128, with results detailed in Table 3. For LLaMA3-70B, LaRoSA achieves notable speed-ups of up to 1.45× and 1.90× at 50% and 75% sparsity, respectively. The speed-up for 0% sparsity is lower than dense model, due to the additional computation introduced by residual adapters and sparsification functions. Nonetheless, we observe overall inference efficiency improvements as sparsity increases. The detailed extra computation in TFLOPS are listed in Table 19.
@@
Impact of Initial Tokens. As noted in Section 5.4.3 of the TEAL paper, TEAL applies activation sparsification to only 99% of tokens during the prefill stage. This approach is attributed to attention sinks (Xiao et al., 2024), which cause more severe degradation when sparsifying the initial tokens. Consequently, TEAL skips the first 1% of tokens to maintain model performance. However, we argue that this issue arises due to a mismatch between the actual and calibrated thresholds for each token. As illustrated in Figure 3(b), this mismatch is particularly pronounced at the beginning of sequences, leading to TEAL’s suboptimal performance on initial tokens. In contrast, LaRoSA and CATS apply sparsification uniformly across all tokens without special treatment. For a fair comparison, we implement full activation sparsification for TEAL, which results in slightly worse performance than originally reported by TEAL. This outcome also underscores the lack of robustness in using empirical thresholds.
@@
Dense 6.13 32.94 (1.00$\times$) 5.47 29.26 (1.00$\times$)
$\text{Baseline}_{25\%}$ 6.29 32.94 (1.00$\times$) 5.53 29.26 (1.00$\times$)
$\text{Baseline}_{25\%}$ + $\mathbf{Q}_{M}$ 6.35 32.94 (1.00$\times$) 5.64 29.26 (1.00$\times$)
$\text{Baseline}_{25\%}$ + $\mathbf{Q}_{B}$ 6.23 37.22 (1.13$\times$) 5.51 33.06 (1.15$\times$)
$\text{Baseline}_{25\%}$ + $\mathbf{Q}_{L}$ 6.23 34.91 (1.06$\times$) 5.51 31.46 (1.07$\times$)
$\text{Baseline}_{50\%}$ 7.71 32.94 (1.00$\times$) 6.02 29.26 (1.00$\times$)
$\text{Baseline}_{50\%}$ + $\mathbf{Q}_{M}$ 7.76 32.94 (1.00$\times$) 6.03 29.26 (1.00$\times$)
$\text{Baseline}_{50\%}$ + $\mathbf{Q}_{B}$ 7.21 37.22 (1.13$\times$) 5.86 33.06 (1.15$\times$)
$\text{Baseline}_{50\%}$ + $\mathbf{Q}_{L}$ 7.22 34.91 (1.06$\times$) 5.87 31.46 (1.07$\times$)
$\text{Baseline}_{60\%}$ 9.39 32.94 (1.00$\times$) 6.91 29.26 (1.00$\times$)
$\text{Baseline}_{60\%}$ + $\mathbf{Q}_{M}$ 9.43 32.94 (1.00$\times$) 6.92 29.26 (1.00$\times$)
$\text{Baseline}_{60\%}$ + $\mathbf{Q}_{B}$ 8.56 37.22 (1.13$\times$) 6.40 33.06(1.15$\times$)
$\text{Baseline}_{60\%}$ + $\mathbf{Q}_{L}$ 8.57 34.91( 1.06$\times$) 6.40 31.46 (1.07$\times$)
@@ 5.3 Ablation Study
Impact of Sparsification Function.We ablate four activation sparsification functions in LaRoSA: 1) TEAL refers to the function that uses a fixed cutoff threshold to prune low-magnitude entries, as described in Equation 5. 2) TopK denotes the approach of introducing sparsity directly using the TopK function. 3) TopK+GS indicates an enhancement of TopK by incorporating Grid Search to obtain optimal coefficients $\alpha$ for different inputs within the same block. 4) LaRoSA represents the complete method, which introduces a layerwise rotation matrix ${\mathbf{Q}}_{l}$ on top of TopK+GS. Table 5 shows that the actual sparsity of the TEAL exhibits fluctuations and fails to achieve a 50% model-level sparsity. In contrast, TopK function can consistently maintain the actual sparsity of each input activation around 50% and significantly reducing PPL. By introducing the optimal sparsity coefficient $\alpha$ found through Grid Search, the PPL loss can be further minimized while keep the overall model sparsity remains unchanged. Finally, incorporating layerwise orthogonal rotation matrices further enhance model’s performance.
Model-level vs. Layer-level vs. Block-level Rotation. We comprehensively investigate the number and sources of rotation matrices and their impact on the performance and efficiency of sparsified models. Three types of rotation matrices are analyzed: 1) Model-level rotation: This approach averages the covariance matrices of input across all layers to obtain a single and shared ${\mathbf{Q}}_{M}$ for all layers, without adding residual adapters. 2) Block-level rotation: Each Attention and MLP block in a layer receives its own rotation matrix ${\mathbf{Q}}_{B}$, derived from its input’s covariance matrix. Thus, a model with $L$ layers generates $2L$ rotation matrices and $2L$ residual adapters. 3) Layer-level rotation: Used by LaRoSA, this strategy assigns a unique rotation matrix ${\mathbf{Q}}_{L}$ to each layer based on its input covariance matrix, resulting in $L$ rotation matrices and $L$ residual adapters. As shown in Table 6, these rotation types result in different additional computation overhead and distinct improvements in PPL. Although ${\mathbf{Q}}_{M}$ does not introduce additional overhead, it actually results in worse performance compared to the baseline. Distinct ${\mathbf{Q}}_{B}$ for each block significantly increases computational overhead, yet offers only marginal advantages over ${\mathbf{Q}}_{L}$. Consequently, the layerwise rotation matrix ${\mathbf{Q}}_{L}$ emerges as the optimal choice, as it increases limited computations with notable performance enhancements.
@@
Influence of Calibration Datasets. We apply LaRoSA to sparsify the LLaMA2-7B and LLaMA3-8B models using different calibration datasets (i.e. WikiText2 and Alpaca (Taori et al., 2023)), with results shown in Table 7. It can be observed that the choice of calibration datasets has a relatively minimal effect on sparsified model performance. This is due to our method utilize calibration data solely for calculating covariance matrices of each layer’s input activation, rather than for empirical distribution computation as employed by methods like CATS and TEAL. Additionally, we validate the performance of LaRoSA within the reasoning model on complex tasks. As shown in Table 8, the covariance matrices of input activations within the same layer exhibit a high degree of cosine similarity across different data inputs. This observation suggests that the acquisition of layerwise rotation matrices is not strongly correlated with the calibration data. The performance in zero-shot and more complex tasks further demonstrates that the calibration data has minimal impact on the model’s performance. This ablation study highlights the robustness of our LaRoSA approach.
Compatiblity with Quantization. We demonstrate that LaRoSA also shows high compatibility with latest SOTA weight quantization methods. In detail, we experimented with LaRoSA using 4-bit quantization(W4A16) on model weights by employing RTN (Dettmers et al., 2022), GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2024b) and OmniQuant (Shao et al., 2024) methods. Table 9 displays the perplexity and zero-shot accuracy results. With 4-bit GPTQ quantization, LaRoSA achieves 64.64% zero-shot accuracy and 5.74 PPL for LLaMA2-7B at 25% sparsity. These results are nearly on par with the the GPTQ quantized model’s 64.89% accuracy and 5.69 PPL. This indicates that the empirical errors from sparsity and quantization are somewhat orthogonal to each other. The compatibility of LaRoSA with weight quantization highlights potential efficiency improvements through optimized kernels that integrate both sparsification and quantization operations.

### v2

@@ 4.2 Layerwise Orthogonal Rotation
The limitations of magnitude pruning drive us to explore a more effective approach for identifying activations that are inherently easier to prune. Inspired by rotation-based quantization methods (ashkboos2024quarot), we attempt to obtain a rotated activation through orthogonal transformations, after which the channel importance can be easily distinguished. We propose using an orthogonal matrix ${\mathbf{Q}}_{l}$ to rotate and prune the input activation (i.e., ${\bm{h}}_{1}$ and ${\bm{h}}_{3}$) of $l$-th layer, and reverse the rotation using $\mathbf{Q}^{\top}_{l}$ after pruning. To construct ${\mathbf{Q}}_{l}$, we use PCA to leverage it’s inherent ability to identify principal components. Specifically, we select a calibration dataset consist of $M$ sequences, feed it into the model and forward pass through each layer. Let ${\mathbf{X}}_{l}^{i}$ denotes the input activation of layer $l$ for $i$-th sequence. We compute the covariance matrix ${\mathbf{X}}_{l}^{i}({\mathbf{X}}_{l}^{i})^{T}$ of each ${\mathbf{X}}_{l}^{i}$ and average over all sequences:
@@
It can be observed that each layer must use the same ${\mathbf{Q}}_{l}$ due to residual connections between layers. We initially attempt to use a single global $\mathbf{Q}$} for all layers, but quickly find out that the optimal orthogonal matrix for each layer tends to be very different. To address this, a residual adapter ${\mathbf{Q}}_{l}^{T}{\mathbf{Q}}_{l+1}$ is applied to the output of each layer to realize layer-specific rotation, as shown in Figure˜2. Although this introduce small additional operations, we demonstrate in Section˜5.3 that such operations preserve the optimal rotation for each layer and improve the overall efficiency. In Appendix A, we provide a theoretical analysis of rotation-based activation pruning, demonstrating that the constructed rotation matrix effectively reduces empirical error within each layer, thereby outperforming magnitude pruning methods in accuracy.
@@ 4.3 Introducing Consistent Activation Sparsity
where ${\mathbf{Q}}_{l}^{T}{\mathbf{W}}_{l}^{T}$ can be precomputed and merged into the weights to avoid online computation. Therefore, Equation˜9 can be rewritten as :
@@ 4.4 Hardware-efficient Customized Kernels
Introducing activation sparsity alone is insufficient to achieve end-to-end LLM inference acceleration on hardware (frantar-sparsegpt). To translate the activation sparsity induced by LaRoSA into wall-clock inference speed-up, we need to implement a fast and efficient customized GPU kernel. Building on the Triton-based kernel introduced by DejaVu (dejavu) and TEAL (liu2024teal), we implement a GEMV kernel that provides practical and consistent token generation speed-up. 1) We store weight matrices $\mathbf{W}$} in column-major format in memory to facilitate the selective loading of weight columns corresponding to non-zero activations, as memory coalescing could significantly improve memory bandwidth utilization (dao2022flashattention; Ivanov2020DataMI). 2) We fuse the Top-K function into the matrix-vector multiplication, identifying the indices of the columns in the weights that need to be retained during the forward pass. 3) We selectively load the sparse activations and corresponding weight columns from memory, transfer them to the computation units, perform the multiplication, and store the result.
@@ 5 Experiments
Table 1: Comparison of the average accuracy on seven zero-shot tasks and 5-shot MMLU. The first column indicates the actual model-level sparsity. CATS cannot achieve 50% model-level sparsity as it only sparsifies the MLP’s Down projections. Full results are in  Appendix˜C.
@@ 5.1 Experimental Setups
Models and Baselines. To test the robustness and scalability of LaRoSA, we conducted experiments using 4 leading LLMs: LLaMA3 (llama3), LLaMA2 (llama2), Mistral (jiang2023mistral), and Qwen2.5 (qwen2). The model size is selected from 7B and 70B parameters. We compare with state-of-the-art baselines, including CATS (lee2024cats) and TEAL (liu2024teal). For SOTA baselines, we choose not to quote the results from their papers in order to make fair comparisons under the same model-level sparsity setting and to impose full sparsity in the input sequence length dimension.
Implementation Details. We use the WikiText2 train set (wikitext) as calibration dataset for LaRoSA and other reproducible works. We randomly select 16 sequences with sequence length of 2048 tokens to compute the rotation matrices $\mathbf{Q}$} for LaRoSA and empirical distributions for CATS and TEAL. The computation of $\mathbf{Q}$} is performed on 8x80G A100 GPUs, taking approximately 12 minutes to complete for the LLaMA3 70B model. For sparsity coefficient $\alpha$, we employ Grid Search to find the optimal hyperparameter for each model, as shown in Appendix˜B Algorithm˜1. The optimal $\alpha$ for each activation type of models is presented in Appendix˜B Table˜11. For a fair comparison, we implement full activation sparsification at prefill stage for all methods including TEAL.
Evaluation. We report the perplexity (PPL) score on WikiText2 test set for generation task. All models are evaluated on the same 128 random samples with a 2048-token context length. For comprehensive assessment, we utilize lm-evaluation-harness(eval-harness) to assess LaRoSA performance on a variety of downstream tasks, including 0-shot tasks for ARC-Easy (clark2018think), ARC Challenge (clark2018think), PIQA (bisk2020piqa), BoolQ (clark2019boolq), HellaSwag (zellers2019hellaswag), OpenBookQA (mihaylov2018can) and WinoGrande (sakaguchi2021winogrande). We also report 5-shot accuracy for MMLU (hendrycks2020measuring).
@@
[…]
@@ 5.2 Main Results
Language Generation Task. First, we evaluate the accuracy of LaRoSA with SOTA baselines on language generation task. As indicated in Table˜2, LaRoSA notably outperforms other baselines on perplexity across various models and different sparsity settings. Under a low sparsity setting like 25%, LaRoSA achieves a minimal PPL loss of merely 0.1 on LLaMA3-70B. This significantly surpasses previous methods, such as CATS with a PPL loss of 0.71 and TEAL with a PPL loss of 1.1. LaRoSA maintains its advantage even at high sparsity levels, such as 60%. For instance, on Qwen2.5-72B, LaRoSA incurs only a 1.03 PPL loss, considerably outperforming TEAL, which suffers a loss of 3.87. Moreover, on LLaMA2-7B, LaRoSA at 50% sparsity even outperforms the previous SOTA methods under a 25% sparsity setting, highlighting its superior efficiency. In TEAL’s paper, they only sparsify 99% of tokens in the prefill stage because the empirical thresholds they use struggle to sparsify initial tokens. In contrast, LaRoSA does not suffer this performance degradation and can achieve full sparsity in the input sequence length dimension.
@@
Complex Problem Solving Tasks. To further evaluate the effectiveness and robustness of LaRoSA, we conduct experiments on complex tasks such as MATH500 (math500), GPQA-Diamond (rein2024gpqa), and AIME’24 (aime) using two reasoning models: DeepSeek-R1-Distill-Llama3-8B and DeepSeek-R1-Distill-Qwen2.5-7B (deepseekr1). We utilize the Alpaca dataset, which is not inherently related to complex reasoning tasks, to derive the rotation matrix ${\mathbf{Q}}_{l}$ for each layer. The evaluation of LaRoSA on these complex tasks is carried out with the Hugging Face Open-R1 repository (openr1). As shown in Table˜4, our experimental results indicate that at a sparsity level of 25%, both reasoning models exhibit only a slight performance degradation on complex tasks. This finding demonstrates that the LaRoSA method maintains its effectiveness even when utilizing data that is not directly related to complex reasoning tasks for generating the rotation matrix.
@@
Inference Speed Up. We demonstrate consistent end-to-end efficiency improvements with LaRoSA. For this, we collected ten samples, each consisting of 128 tokens, from various test datasets and generated new sequences with lengths ranging from 128 to 2048 tokens. This was done to evaluate performance across different generation lengths. We applied 50% sparsity to LLaMA3-8B and Qwen2.5-7B using optimal sparse patterns obtained from LaRoSA and TEAL. As depicted in Figure˜4, our LaRoSA realizes more consistent wall-clock time speed-ups compared to TEAL. This improvement is attributed to the TopK strategy which precisely zeros out activation entries needed for a specific sparsity level. As the output sequence length increases, LaRoSA outperforms TEAL in overall speed-up efficiency, demonstrating its effectiveness and robustness. Moreover, we evaluated the end-to-end token generation speed-up on different types of models at a fixed output sequence length of 128, with results detailed in Table˜3. For LLaMA3-70B, LaRoSA achieves notable speed-ups of up to 1.45× and 1.90× at 50% and 75% sparsity, respectively. The speed-up for 0% sparsity is lower than dense model, due to the additional computation introduced by residual adapters and sparsification functions. Nonetheless, we observe overall inference efficiency improvements as sparsity increases. The detailed extra computation in TFLOPS are listed in Table˜19.
@@
Impact of Initial Tokens. As noted in Section 5.4.3 of the TEAL paper, TEAL applies activation sparsification to only 99% of tokens during the prefill stage. This approach is attributed to attention sinks (xiao2024efficient), which cause more severe degradation when sparsifying the initial tokens. Consequently, TEAL skips the first 1% of tokens to maintain model performance. However, we argue that this issue arises due to a mismatch between the actual and calibrated thresholds for each token. As illustrated in Figure˜3(b), this mismatch is particularly pronounced at the beginning of sequences, leading to TEAL’s suboptimal performance on initial tokens. In contrast, LaRoSA and CATS apply sparsification uniformly across all tokens without special treatment. For a fair comparison, we implement full activation sparsification for TEAL, which results in slightly worse performance than originally reported by TEAL. This outcome also underscores the lack of robustness in using empirical thresholds.
@@
[…]
@@ 5.3 Ablation Study
Impact of Sparsification Function.We ablate four activation sparsification functions in LaRoSA: 1) TEAL refers to the function that uses a fixed cutoff threshold to prune low-magnitude entries, as described in Equation˜5. 2) TopK denotes the approach of introducing sparsity directly using the TopK function. 3) TopK+GS indicates an enhancement of TopK by incorporating Grid Search to obtain optimal coefficients $\alpha$ for different inputs within the same block. 4) LaRoSA represents the complete method, which introduces a layerwise rotation matrix ${\mathbf{Q}}_{l}$ on top of TopK+GS. Table˜5 shows that the actual sparsity of the TEAL exhibits fluctuations and fails to achieve a 50% model-level sparsity. In contrast, TopK function can consistently maintain the actual sparsity of each input activation around 50% and significantly reducing PPL. By introducing the optimal sparsity coefficient $\alpha$ found through Grid Search, the PPL loss can be further minimized while keep the overall model sparsity remains unchanged. Finally, incorporating layerwise orthogonal rotation matrices further enhance model’s performance.
Model-level vs. Layer-level vs. Block-level Rotation. We comprehensively investigate the number and sources of rotation matrices and their impact on the performance and efficiency of sparsified models. Three types of rotation matrices are analyzed: 1) Model-level rotation: This approach averages the covariance matrices of input across all layers to obtain a single and shared ${\mathbf{Q}}_{M}$ for all layers, without adding residual adapters. 2) Block-level rotation: Each Attention and MLP block in a layer receives its own rotation matrix ${\mathbf{Q}}_{B}$, derived from its input’s covariance matrix. Thus, a model with $L$ layers generates $2L$ rotation matrices and $2L$ residual adapters. 3) Layer-level rotation: Used by LaRoSA, this strategy assigns a unique rotation matrix ${\mathbf{Q}}_{L}$ to each layer based on its input covariance matrix, resulting in $L$ rotation matrices and $L$ residual adapters. As shown in Table˜6, these rotation types result in different additional computation overhead and distinct improvements in PPL. Although ${\mathbf{Q}}_{M}$ does not introduce additional overhead, it actually results in worse performance compared to the baseline. Distinct ${\mathbf{Q}}_{B}$ for each block significantly increases computational overhead, yet offers only marginal advantages over ${\mathbf{Q}}_{L}$. Consequently, the layerwise rotation matrix ${\mathbf{Q}}_{L}$ emerges as the optimal choice, as it increases limited computations with notable performance enhancements.
@@
Influence of Calibration Datasets. We apply LaRoSA to sparsify the LLaMA2-7B and LLaMA3-8B models using different calibration datasets (i.e. WikiText2 and Alpaca (alpaca)), with results shown in Table˜7. It can be observed that the choice of calibration datasets has a relatively minimal effect on sparsified model performance. This is due to our method utilize calibration data solely for calculating covariance matrices of each layer’s input activation, rather than for empirical distribution computation as employed by methods like CATS and TEAL. Additionally, we validate the performance of LaRoSA within the reasoning model on complex tasks. As shown in Table 8, the covariance matrices of input activations within the same layer exhibit a high degree of cosine similarity across different data inputs. This observation suggests that the acquisition of layerwise rotation matrices is not strongly correlated with the calibration data. The performance in zero-shot and more complex tasks further demonstrates that the calibration data has minimal impact on the model’s performance. This ablation study highlights the robustness of our LaRoSA approach.
Compatiblity with Quantization. We demonstrate that LaRoSA also shows high compatibility with latest SOTA weight quantization methods. In detail, we experimented with LaRoSA using 4-bit quantization(W4A16) on model weights by employing RTN (llm_int8), GPTQ (frantar-gptq), AWQ (lin2023awq) and OmniQuant (shao2024omniquant) methods. Table˜9 displays the perplexity and zero-shot accuracy results. With 4-bit GPTQ quantization, LaRoSA achieves 64.64% zero-shot accuracy and 5.74 PPL for LLaMA2-7B at 25% sparsity. These results are nearly on par with the the GPTQ quantized model’s 64.89% accuracy and 5.69 PPL. This indicates that the empirical errors from sparsity and quantization are somewhat orthogonal to each other. The compatibility of LaRoSA with weight quantization highlights potential efficiency improvements through optimized kernels that integrate both sparsification and quantization operations.